OpenStudy (anonymous):

how do you solve \sqrt{3k-11}=\sqrt{5-k}

OpenStudy (anonymous):

Square it

OpenStudy (anonymous):

i tried still dont get it because i think im missing something but dont know what it is

OpenStudy (anonymous):

square both sides...so u would have; 3k-11=5-k; k=4

OpenStudy (anonymous):

thank u